OpenStudy (anonymous):

Evaluate the indefinite integral. integral((3dx)/(xln(2x))) Integral = ________ +C

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{3 dx}{x \ln(2x)}\]

OpenStudy (anonymous):

I'll give you a hint, do a u-substitution with u=ln(2x)

OpenStudy (anonymous):

d/dx of ln(2x)=2*1/(2x)=1/x

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{3}{x \ln(x)} dx\]

OpenStudy (anonymous):

Your "du" calculation is correct, but the substitution isn't, the integral would change to this. Make sense?